OpenStudy (anonymous):

need help on this one −x+5y=14;(−5,−2) Write the equation of a line that is perpendicular to the given line and that passes through the given point.

OpenStudy (anonymous):

A. y= 5x - 27 B. y= -1/5x - 27 C. y= -1/5 x - 15 \ D. y= -5x - 27 these are my choices

OpenStudy (anonymous):

@Philippians413 can you give me a quick answer on this one

OpenStudy (anonymous):

oh umm hold on i will try no promises though after that last one i was wrong on i don't how it will turn out but I will try hold on a sec

OpenStudy (anonymous):

ok thank you try your best any effort is good enough for me

OpenStudy (anonymous):

I am so sorry I have been on the computer all day and my brain is procrastinating i am trying as fast as I can

OpenStudy (anonymous):

its ok i got plenty of time..

OpenStudy (anonymous):

I am pretty sure it is a I found a graph calculator and a is perpendicular to -x+5y=14

OpenStudy (anonymous):

hope that helps and is right this time lol

OpenStudy (anonymous):

actually that's no in one of my choices as an answer.... A. y= 5x - 27 B. y= -1/5x - 27 C. y= -1/5 x - 15 \ D. y= -5x - 27 these are my choices

OpenStudy (anonymous):

no A is the answer y=5x-27 is the answer to what is perpendicular to -x+5y=14

OpenStudy (anonymous):

ohhhhh ok thank you :) lol

OpenStudy (anonymous):

lol no prob